# Distributional Vision-Language Alignment by Cauchy-Schwarz Divergence

### v1

@@ C.1 T2I details
We use Kandinsky v2.2, an unCLIP-type model that utilizes CLIP ViT-bigG-14-laion2B-39B-b160k with 1280 projection dimensions for text and image encoders. Kandinsky v2.2 employs a latent diffusion model and MOVQ (Zheng et al., 2022) as the decoder to generate images of size $512\times 512$ from the given image representation. When using the Kandinsky decoder, we apply $50$ denoising steps (Ho et al., 2020) with a classifier-free guidance scale of 7.5 (Ho & Salimans, 2022).
@@
To ensure a fair comparison, our adapter module has the same architecture as Eclipse (Patel et al., 2024), which is based on the standard PriorTransformer model (Ramesh et al., 2022) but modified to be time-independent. Specifically, it consists of 10 layers with 16 attention heads, each having a head dimension of 32. The embedding dimension is 768/1280, with three additional embeddings. The model does not use time embeddings and has a dropout rate of 0.0.
@@
Appendix D More experimental results
@@
D.1 More visualization
D.2 More visualization for token alignment

### v2

@@ F.1 T2I details
We use Kandinsky v2.2, an unCLIP-type model that utilizes CLIP ViT-bigG-14-laion2B-39B-b160k with 1280 projection dimensions for text and image encoders. Kandinsky v2.2 employs a latent diffusion model and MOVQ [57] as the decoder to generate images of size $512\times 512$ from the given image representation. When using the Kandinsky decoder, we apply $50$ denoising steps [58] with a classifier-free guidance scale of 7.5 [59].
@@
To ensure a fair comparison, our adapter module has the same architecture as Eclipse [15], which is based on the standard PriorTransformer model [3] but modified to be time-independent. Specifically, it consists of 10 layers with 16 attention heads, each having a head dimension of 32. The embedding dimension is 768/1280, with three additional embeddings. The model does not use time embeddings and has a dropout rate of 0.0. For the text to image generation task, in order to use the pretrained image generator, we only use the text adapter. For the retrieval and classification, we use adapters for both modalities.
@@
Appendix G More experimental results
Image-text classification.
We compare with CLIP-Adapter [13] on the image classification task following their few-shot classification setting. We fine-tune the adapter based on ViT-B/16 with 16-shots subset for each of the 11 datasets. The results are provided in the following table. With better alignment, our method consistently performs better.
Table 5: Comparison with CLIP-Adapter on the image classification task. Our methods performs consistently better on various datasets.
Method ImageNet Caltech101 DTD EuroSAT FGVCAircraft Food101 Flowers102 OxfordPets StanfordCars SUN397 UCF101 Average
[13] 71.1 94.4 70.9 85.7 42.8 83.2 96.0 92.1 78.6 75.0 82.8 79.3
Ours 72.9 95.0 72.3 87.2 44.4 85.8 97.5 93.0 81.9 76.2 84.0 80.9
Hyperparameter Sensitivity Analysis. We conducted a sensitivity analysis on the two key hyperparameters, $\lambda$ (the weight for InfoNCE) and $\sigma$ (the Gaussian kernel width). For efficiency, we evaluated on a subset of 10 000 MSCOCO training samples and report Fréchet Inception Distance (FID) as the metric.
Table 6 shows that our method is robust to moderate variations in both $\lambda$ and $\sigma$, with only minor FID fluctuations over a wide range.
Table 6: Sensitivity of FID to $\lambda$ and $\sigma$.
$\lambda$ 0.01 0.1 1 10
FID 81.34 32.51 29.86 65.79
(a) $\lambda$ sensitivity
$\sigma$ 0.1 0.5 1 1.5
FID 30.58 27.79 29.86 31.49
(b) $\sigma$ sensitivity